OpenStudy (anonymous):

Hunus: Last one! :) Jean wants to build a fence to enclose an area for a new rose garden. She can afford at most 150 feet of fencing. The length of the rectangular garden will be 5 feet more than the width, w. Which inequality best describes the possible width of her garden?

OpenStudy (hunus):

\[l=w+5\] \[2l+2w \le 150\] \[2(w+5) +2w \le 150\] \[2w +10 +2w \le 150\] \[4w \le 140\] \[w \le 35\]

OpenStudy (anonymous):

a\[a. w+(w+5)\le150 b. 2w+2(w+5)\le150 c. 2w+2(w+5)\ge150 d. w+(w+5)\ge150\]

OpenStudy (hunus):

b. 2w+2(w+5)≤150

OpenStudy (anonymous):

thank you for your patience!! Are you an algebra expert??

OpenStudy (hunus):

Something along those lines lol

OpenStudy (anonymous):

haha, what about geometry?

OpenStudy (hunus):

Is it hard geometry? lol

OpenStudy (anonymous):

I just started so nope :P

OpenStudy (anonymous):

can I post one?

OpenStudy (hunus):

sure